OpenStudy (anonymous):

hiw would i solve x/1.8=72

OpenStudy (anonymous):

x/1.8=72 x/1.8*1.8=72*1.8 x=129.6

OpenStudy (anonymous):

Multiply both sides by 1.8.

OpenStudy (anonymous):

multiply both sides with 1.8

OpenStudy (anonymous):

cross multiply x = 1.8 * 72 = 129.6

OpenStudy (anonymous):

ty